OpenStudy (anonymous):

. A jet ski company charges a flat fee of $24.00 plus $2.25 per hour to rent a jet ski. Another company charges a fee of $23.00 plus $2.50 per hour to rent the same jet ski. Using a graphing calculator, find the number of hours for which the costs are the same. Round your answer to the nearest whole hour. (1 point) A• 9 B• 8 C• 4 D•3

OpenStudy (anonymous):

Can somebody HELP ME PLEASE

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

somebody can

OpenStudy (anonymous):

its just not going to be me sucker

OpenStudy (thejax):

Use the sentences to write 2 equations, with the variables being: t = number of hours C = total cost of rental First jet ski company: C = 2.25t + 24 Second jet ski company: C = 2.5t + 23 Now enter these two equations into your graphing calculator and find the point at which they intersect. Since Yahoo Answers is not a graphing calculator, you can figure out where they intersect if you set the cost of the first jet ski company equal to the cost of the second jet ski company. 2.25t + 24 = 2.5t + 23 Subtract 2.25t from each side 24 = .25t + 23 Subtract 23 from each side 1 = .25t Divide both sides by .25 t = 4 Therefore the cost is the same at 4 hours.

OpenStudy (austinl):

You have two lines. \(y=2.25x+24\) \(y=2.50x+23\) Graph them and see where they intersect!

OpenStudy (austinl):

@TheJax It said using a graphing calculator..

OpenStudy (anonymous):

@austinL so I want to say (A)

OpenStudy (austinl):

Do you have a TI-83 or 84?

OpenStudy (anonymous):

84 @austinL

OpenStudy (austinl):

Ok, perfect. Press Y= Then type in the two lines I listed above. Press ZOOM, then choose 0/FIT Press 2ND>TRACE Then 5/intersect Choose one line as 1st curve Choose the other as second line (it should automatically change, you can just press enter twice) Press enter through "Guess?" That gives you the intersection point.

OpenStudy (anonymous):

SO the answer will be (c)

OpenStudy (austinl):

Correct. @TheJax 's way is correct, but I wanted to show you how to do it according to the instructions. :)

OpenStudy (anonymous):

ok thanks so much

OpenStudy (austinl):

No problem!

OpenStudy (thejax):

np, i feel u deserve it! @austinL

OpenStudy (thejax):

Thanks for giving a good competition to a Engg Grad!`